OpenStudy (anonymous):

Could someone please help quickly? Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A. (0, 6) B. (–5, –2) C. (6, 0) D. (–2, –5)

jimthompson5910 (jim_thompson5910):

let's test choice A

jimthompson5910 (jim_thompson5910):

(0,6) is in the form (a,b) a = 0 b = 6

jimthompson5910 (jim_thompson5910):

4a – 5b = 24 4*0 – 5*6 = 24 ... plug in a = 0, b = 6 0 - 30 = 24 -30 = 24 ... false

jimthompson5910 (jim_thompson5910):

since the result is false, this means (0,6) is NOT a solution to 4a – 5b = 24

jimthompson5910 (jim_thompson5910):

you are doing the same for the other answer choices

OpenStudy (anonymous):

...

jimthompson5910 (jim_thompson5910):

the most descriptive response I've ever seen

OpenStudy (anonymous):

I see.